OpenStudy (anonymous):

Solve. x divided by eight equals five divided by forty x equals one x equals five-eighths x equals one and three-fifths x equals five

OpenStudy (anonymous):

x/8 = 5/40 x = (8 * 5)/40 = 1

OpenStudy (anonymous):

THANKS